# Hierarchical Lorentz Mirror Model: Normal Transport and a Universal $2/3$ Mean–Variance Law

### v1

@@ Abstract
The Lorentz mirror model provides a clean setting to study macroscopic transport generated solely by quenched environmental randomness. We introduce a hierarchical version that admits an exact recursion for the distribution of left–right crossings, and prove normal transport: the mean conductance scales as (cross-section)/(length) for all length scales if $d\geq 3$. A Gaussian approximation, supported by numerics, predicts that, in the marginal case $d=2$, this scaling acquires a logarithmic correction and that the variance-to-mean ratio of conductance converges to the universal value $2/3$ (the “$2/3$ law”) for all $d\geq 2$. We conjecture that both effects persist beyond the hierarchical setting. We finally provide numerical evidence for the $2/3$ law in the original Lorentz mirror model in $d=3$, and interpret it as a universal signature of normal transport induced by random current matching.
@@ A YouTube video discussing the background and the main results of the paper is available:
Deriving macroscopic normal transport—as described by diffusive laws such as Fick’s, Ohm’s, and Fourier’s laws—from microscopic deterministic dynamics remains a central challenge in nonequilibrium statistical mechanics BonettoLebowitzReyBellet ; LepriLiviPoliti ; DharReview ; LebowitzSpohn . Lorentz-type models, in which particles move deterministically through an externally imposed (often disordered) environment, provide a natural testing ground for this program Lorentz1905 ; DettmannReview ; vanBeijeren . The Lorentz mirror model introduced by Ruijgrok and Cohen Ruijgrok is a particularly clean lattice realization: local scattering is deterministic, so any normal transport must be generated by quenched environmental randomness rather than by stochastic forcing or chaotic microscopic dynamics Ruijgrok ; Basile ; Sanders ; BunimovichTroubetzkoy ; BunimovichTumarkin ; CohenWang ; Lefevere2015 ; ChiffaudelLefevere ; ElboimGloriaHernandez ; Lefevere2025 ; Grimmett ; KozmaSidoravicius ; Ryan2021 ; GrimmettSaintFlour .
In the present Letter, we propose and study a hierarchical version of the Lorentz mirror model. We rigorously establish normal transport in dimensions $d\geq 3$, and conjecture weakly anomalous transport with a logarithmic correction in the marginal case $d=2$. Most strikingly, we conjecture that the ratio of the sample-to-sample variance of the conductance to its mean approaches the universal value $2/3$ not only in the hierarchical model but also in the original mirror model, suggesting a novel universal signature of normal transport induced by random current matching. A video discussing the background and the main results of the paper is available video .
Original Lorentz mirror model.— Let us describe the Lorentz mirror model, starting with $d=2$. Consider the $L\times L$ square lattice with periodic boundary conditions in the vertical direction and open boundary conditions in the horizontal direction. We attach one external edge to each boundary vertex on the left and on the right (thick lines in Fig. 1), so that every vertex has four incident edges. (This particular geometry, suited to the study of transport properties, was first introduced in Lefevere2015 .) At each vertex, we independently choose one of the three pairings of the four incident edges, each with probability $1/3$. These local pairings decompose the edges into a collection of disjoint trajectories and induce a global perfect matching of the external edges; see Fig. 1. As the name suggests, in $d=2$ the local pairing may be viewed as arising from reflections by randomly placed mirrors SM .
@@
Similarly, for $d\geq 3$, we consider the $d$-dimensional $L\times\cdots\times L$ hypercubic lattice with open boundary conditions in one (horizontal) direction and periodic boundary conditions in the remaining d1d-1 directions. We attach $L^{d-1}$ external edges to each of the two open boundaries, and at each vertex independently choose one of the (2d1)!!(2d-1)!! pairings of the $2d$ incident edges with equal probability. (Note that this does not literally correspond to mirror configurations.) Again, the local wiring induces a global perfect matching of the external edges.
@@
if $\ell\in S(\ell_{\mathrm{L}},\ell_{\mathrm{R}})$, and $K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})=0$ otherwise, where $S(a,b):=2\mathbb{Z}\cap[0,\min\{a,b\}]$. It is normalized as $\sum_{\ell\in S(\ell_{\mathrm{L}},\ell_{\mathrm{R}})}K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})=1$. This follows by observing that the uniform random matching of all $2A_{n}$ interface edges induces, by permutation symmetry, a uniform random matching of the $\ell_{\mathrm{L}}+\ell_{\mathrm{R}}$ distinguished endpoints, and then counting the matchings with exactly $\ell$ left–right pairs. See below for details.
@@
Gaussian heuristic and the $2/3$ law.— A useful (and surprisingly accurate) heuristic is obtained by a Gaussian closure: (i) assume that, for sufficiently large $n$, the probability distribution $P_{n}$ is well approximated by a Gaussian with mean $\mu_{n}=\langle\hat{\ell}\rangle_{n}$ and variance $v_{n}:=\langle(\hat{\ell}-\mu_{n})^{2}\rangle_{n}$; (ii) approximate the interface kernel $K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})$ by a Gaussian in $\ell$,
@@
with the conditional mean $\tilde{\mu}(\ell_{\mathrm{L}},\ell_{\mathrm{R}})$ and conditional variance $\tilde{v}(\ell_{\mathrm{L}},\ell_{\mathrm{R}})$ given above. (In fact, (7) can be proved in an appropriate asymptotic regime.)
@@
Here, the two contributions to $v_{n}$ come from the conditional variance in (7) and from the fluctuations of $\ell_{\mathrm{L}}$ and $\ell_{\mathrm{R}}$, respectively SM .
@@
which is independent of $d$. Therefore $v_{n}/\mu_{n}\to 2/3$ as $n\to\infty$ (the “$2/3$ law”).
In the marginal case $d=2$, the leading drift in (8) cancels; retaining the next-order terms predicts a logarithmic growth
@@
while still $v_{n}/\mu_{n}\to 2/3$; see SM .
Numerical results for the hierarchical model.— We numerically iterated the recursion (2) in $d=3$ and $d=2$ to test the Gaussian-closure predictions. The agreement is excellent. In $d=3$, after only a few iterations, the mean conductance $\mu_{n}$ is essentially proportional to $L_{n}$, and the ratio $v_{n}/\mu_{n}$ becomes numerically indistinguishable from $2/3$. In the marginal case $d=2$, after reducing discreteness effects by starting from a moderately large initial crossing number $A_{0}$, we find that $\mu_{n}$ grows linearly in $n=\log_{2}L_{n}$ with slope consistent with $1/12$, and again $v_{n}/\mu_{n}$ rapidly locks onto $2/3$. We also observe that the standardized law approaches a Gaussian (in Kolmogorov distance) rapidly in $d=3$, apparently exponentially in $n$, whereas in $d=2$ the convergence is much slower and compatible with a $n^{-1}$ decay. See SM for details.
The remarkably fast stabilization of $v_{n}/\mu_{n}$ to $2/3$ suggests viewing this ratio as a universal amplitude ratio, characterizing a universality class that contains the present hierarchical Lorentz mirror model. As we shall see below, this expectation is justified, in a stronger sense than we initially anticipated.
@@
Numerical results for the original Lorentz mirror model in d=3d=3.— Normal-transport scaling (1) in the original Lorentz mirror model for $d=3$ was first supported numerically in ChiffaudelLefevere . The proportionality constant (conductivity) $\kappa:=\lim_{L\uparrow\infty}\bar{\mu}(L)/L$ was derived in Lefevere2025 from a multiscale analysis under a closure hypothesis on trajectory correlations. This constant is nonuniversal and depends on the microscopic pairing rule.
@@
Discussion.— We introduced a hierarchical Lorentz mirror model that captures deterministic transport generated solely by quenched environmental randomness through random matching. For this model, we proved that the conductance exhibits normal-transport scaling, even though the microscopic dynamics is far from diffusive and many trajectories are closed, as in the original mirror model. More precisely, we established that $\mu_{n}\propto A_{n}/L_{n}$ in all dimensions $d\geq 3$. The marginal case $d=2$ remains open: the Gaussian approximation and numerics suggest that $\mu_{n}$ grows logarithmically with the system size.
A second outcome is the “$2/3$ law”: whenever $\mu_{n}$ grows with scale, the variance satisfies $v_{n}/\mu_{n}\to 2/3$. In the hierarchical model, this emerges from the Gaussian closure leading to the recursion (9), and is supported numerically. Moreover, we provide numerical evidence that the same ratio holds for the original Lorentz mirror model in $d=3$, and appears robust under variations of the local pairing rule. This motivates the conjecture that the “$2/3$ law” is universal within a broad class of models.
Our results suggest that $d=2$ is the unique marginal dimension of the hierarchical model, and we conjecture the same for the original model. A single injected trajectory may be viewed as a non-backtracking walk with strong memory (encoding the quenched random environment), suggesting a comparison with “true” (myopic) self-avoiding walks and related self-repelling processes, for which $d=2$ is also marginal and which become diffusive in $d\geq 3$ AmitParisiPeliti ; Toth1995 ; TothWerner1998 ; HorvathTothVeto2012 . At the same time, if crossings were produced by essentially independent diffusive trajectories, one would expect Poisson counting statistics, for which $v/\mu$ equals $1$ rather than approaching the sub-Poisson value $2/3$. This contrast suggests that the global current matching—rather than the scaling of an individual trajectory—is the mechanism behind normal transport and the $2/3$ law. A rigorous derivation of the $2/3$ law, and control of the marginal case $d=2$, remain important open problems.
@@ Acknowledgements.
R.L. thanks Gakushuin University for its hospitality during his stay in Tokyo, when the main part of this work was carried out. He also acknowledges the financial support from FJ-LMI IRL 2025 CNRS-The University of Tokyo, which made the visit possible. The present research is supported by JSPS Grants-in-Aid for Scientific Research Nos. 22K03474 (R.L. and H.T.) and 25K07171 (H.T.). We made extensive use of ChatGPT (versions 5.1 and 5.2, Plus/Pro) throughout this project for code development, assistance with proofs and calculations, reference searches, and drafting of the manuscript. In particular, the role of the identity (16) in the proof of Theorem End Matter was identified with the assistance of ChatGPT 5.2 Pro (although, interestingly, it never realized (16) until we pointed out that the quantity could be evaluated).
@@
$u_{n}=\sum_{\ell_{1},\ldots,\ell_{m}}\Bigl(1+\sum_{j=1}^{m}\ell_{j}\Bigr)^{-1}\prod_{j=1}^{m}P_{n}(\ell_{j}),$ (SM.16)
@@
$u_{n}=\sum_{y_{1},\ldots,y_{m}}\Bigl(\sum_{j=1}^{m}y_{j}^{-1}-(m-1)\Bigr)^{-1}\prod_{j=1}^{m}Q_{n}(y_{j}).$ (SM.17)
@@
$(y_{1}^{-1}+c)^{-1}=\frac{y_{1}}{1+cy_{1}}$ (SM.18)
is concave in $y_{1}\in(0,1]$, since its second derivative equals $2c(1+cy_{1})^{-3}\leq 0$. Applying Jensen’s inequality to the $y_{1}$-average in (SM.17) yields
$u_{n}\leq\sum_{y_{2},\ldots,y_{m}}\Bigl(\eta_{n}^{-1}+\sum_{j=2}^{m}y_{j}^{-1}-(m-1)\Bigr)^{-1}\prod_{j=2}^{m}Q_{n}(y_{j}).$ (SM.19)
@@
$u_{n}\leq\bigl(m\eta_{n}^{-1}-(m-1)\bigr)^{-1}=\frac{\eta_{n}}{m-(m-1)\eta_{n}}=\frac{\eta_{n}}{2^{d-1}-(2^{d-1}-1)\eta_{n}}.$ (SM.20)
@@
Substituting (SM.20) (with $n\mapsto n-1$) into (24) and using that $x\mapsto 2x-x^{2}$ is increasing on $[0,1]$, we obtain
$\eta_{n}\leq\frac{2m\,\eta_{n-1}-(2m-1)\eta_{n-1}^{2}}{\bigl(m-(m-1)\eta_{n-1}\bigr)^{2}}.$ (SM.21)
@@
$\qquad=-\frac{\eta^{2}\bigl(\eta(m-1)^{2}-m(m-2)\bigr)^{2}}{m^{2}\bigl(m-(m-1)\eta\bigr)^{2}}\leq 0.$ (SM.22)
Combining (SM.21) and (SM.22) yields the quadratic recursion
$\eta_{n}\leq\frac{2}{m}\eta_{n-1}+\Bigl(\frac{m-1}{m}\Bigr)^{2}\eta_{n-1}^{2}=2^{2-d}\eta_{n-1}+(1-2^{1-d})^{2}\eta_{n-1}^{2}.$ (SM.23)
@@
$x_{n}:=\alpha^{-n}\eta_{n}=\Bigl(\frac{m}{2}\Bigr)^{n}\eta_{n}=2^{(d-2)n}\eta_{n}.$ (SM.24)
Then (SM.23) implies
$x_{n}\leq x_{n-1}+\beta\,\alpha^{\,n-2}x_{n-1}^{2},\qquad\beta:=\Bigl(\frac{m-1}{m}\Bigr)^{2},$ (SM.25)
@@
$\frac{1}{x_{n}}\geq\frac{1}{x_{n-1}}-\beta\,\alpha^{\,n-2},$ (SM.26)
@@
$\frac{1}{x_{n}}\geq(A_{0}+1)-\beta\sum_{j=1}^{n}\alpha^{j-2}\geq(A_{0}+1)-\frac{\beta}{\alpha(1-\alpha)}=(A_{0}+1)-\frac{(m-1)^{2}}{2(m-2)}.$ (SM.27)
@@
$\eta_{n}^{-1}=\alpha^{-n}x_{n}^{-1}\geq\Bigl(A_{0}+1-\frac{(m-1)^{2}}{2(m-2)}\Bigr)\Bigl(\frac{m}{2}\Bigr)^{n}=\Bigl(A_{0}+1-\frac{(2^{d-1}-1)^{2}}{2(2^{d-1}-2)}\Bigr)2^{(d-2)n}.$ (SM.28)
@@
$\mu_{n}\geq\Biggl[1-\frac{1}{A_{0}}\biggl(\frac{(2^{d-1}-1)^{2}}{2(2^{d-1}-2)}-1\biggr)\Biggr]\frac{A_{n}}{L_{n}}-1,$ (SM.29)
@@
The present section provides additional information on our numerical work. We focus on the hierarchical Lorentz mirror model, where the recursion (2) can be iterated essentially exactly (i.e., without Monte Carlo sampling), and we briefly discuss the original (non-hierarchical) Lorentz mirror model in the final subsection.
@@
$P_{n-1}^{*m}(\ell^{\prime}):=\mathop{\sum_{\ell_{1},\ldots,\ell_{m}}}_{(\sum_{j=1}^{m}\ell_{j}=\ell^{\prime})}\prod_{j=1}^{m}P_{n-1}(\ell_{j}),$ (SM.30)
@@
$P_{n}(\ell)=\sum_{\ell_{\mathrm{L}},\ell_{\mathrm{R}}}K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})\,P_{n-1}^{*m}(\ell_{\mathrm{L}})\,P_{n-1}^{*m}(\ell_{\mathrm{R}}).$ (SM.31)
@@
$\mu_{n}\simeq a+b\,n\quad\text{(as $n\uparrow\infty$)},\qquad b\simeq 0.08346,$ (SM.32)

### v2

@@ Abstract
The Lorentz mirror model provides a clean setting to study macroscopic transport generated solely by quenched environmental randomness. We introduce a hierarchical version whose distribution of left–right crossings satisfies an exact recursion. In dimensions $d\geq 3$, we prove normal transport: the mean conductance scales as (cross-section)/(length) on all length scales. A Gaussian closure, supported by numerics, predicts that the variance-to-mean ratio of the conductance converges to the universal value $2/3$ for all $d\geq 2$ (the “$2/3$ law”). We provide numerical evidence for the $2/3$ law in the original (non-hierarchical) Lorentz mirror model in $d=3$, and conjecture that it is a universal signature of normal transport induced by random current matching. In the marginal case $d=2$, our hierarchical recursion reproduces the known scaling of the mean conductance and its variance.
@@ A YouTube video discussing the background and the main results of the paper is available:
Deriving macroscopic normal transport—as described by diffusive laws such as Fick’s, Ohm’s, and Fourier’s laws—from microscopic deterministic dynamics remains a central challenge in nonequilibrium statistical mechanics BonettoLebowitzReyBellet ; LepriLiviPoliti ; DharReview ; LebowitzSpohn . Lorentz-type models, in which particles move deterministically through an externally imposed (often disordered) environment, provide a natural testing ground for this program Lorentz1905 ; DettmannReview ; vanBeijeren . The Lorentz mirror model introduced by Ruijgrok and Cohen RuijgrokCohen is a particularly clean lattice realization: local scattering is deterministic, so any normal transport must be generated by quenched environmental randomness rather than by stochastic forcing or chaotic microscopic dynamics RuijgrokCohen ; ZiffKongCohen1991 ; Basile ; KraemerSanders ; BunimovichTroubetzkoy ; BunimovichTumarkin ; NienhuisRietman ; CohenWang ; OwczarekPrellberg ; BezuidenhoutGrimmett1999 ; Lefevere2015 ; ChiffaudelLefevere ; ElboimGloriaHernandez ; Lefevere2025 ; Grimmett ; KozmaSidoravicius ; Ryan2021 ; GrimmettKesten ; NahumSernaSomozaOrtuno2013 . See, e.g., NahumSernaSomozaOrtuno2013 and references therein for its close relation to loop models.
In the present Letter, we propose and study a hierarchical version of the Lorentz mirror model. We rigorously establish normal transport in dimensions $d\geq 3$, and observe weakly anomalous transport with a logarithmic correction in the marginal case $d=2$. Most strikingly, we conjecture that the ratio of the sample-to-sample variance of the conductance to its mean approaches the universal value $2/3$ not only in the hierarchical model but also in the original mirror model, suggesting a novel universal signature of normal transport induced by random current matching. A video discussing the background and the main results of the paper is available video .
Original Lorentz mirror model.— Let us describe the Lorentz mirror model, starting with $d=2$. The model may be viewed a special case of the compactly packed loop model with crossings NahumSernaSomozaOrtuno2013 . Consider the $L\times L$ square lattice with periodic boundary conditions in the vertical direction and open boundary conditions in the horizontal direction. We attach one external edge to each boundary vertex on the left and on the right (thick lines in Fig. 1), so that every vertex has four incident edges. At each vertex, we independently choose one of the three pairings of the four incident edges, each with probability $1/3$. These local pairings decompose the edges into a collection of disjoint trajectories and induce a global perfect matching of the external edges; see Fig. 1. As the name suggests, in $d=2$ the local pairing may be viewed as arising from reflections by randomly placed mirrors SM .
@@
Similarly, for $d\geq 3$, we consider the $d$-dimensional $L\times\cdots\times L$ hypercubic lattice with open boundary conditions in one (horizontal) direction and periodic boundary conditions in the remaining d1d-1 directions, and attach $L^{d-1}$ external edges to each of the two open boundaries. (This particular geometry, suited to the study of transport properties, was studied in Lefevere2015 .) We then independently choose one of the (2d1)!!(2d-1)!! pairings of the $2d$ incident edges with equal probability at each vertex. (Note that this does not literally correspond to mirror configurations.) Again, the local wiring induces a global perfect matching of the external edges.
@@
if $\ell\in S(\ell_{\mathrm{L}},\ell_{\mathrm{R}})$, and $K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})=0$ otherwise, where $S(a,b):=2\mathbb{Z}\cap[0,\min\{a,b\}]$. It is normalized as $\sum_{\ell\in S(\ell_{\mathrm{L}},\ell_{\mathrm{R}})}K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})=1$. This follows by observing that the uniform random matching of all $2A_{n}$ interface edges induces, by permutation symmetry, a uniform random matching of the $\ell_{\mathrm{L}}+\ell_{\mathrm{R}}$ distinguished endpoints.
@@
Gaussian heuristic and the $2/3$ law.— A useful (and surprisingly accurate) heuristic is obtained by a Gaussian closure SM : (i) assume that, for sufficiently large $n$, the probability distribution $P_{n}$ is well approximated by a Gaussian with mean $\mu_{n}=\langle\hat{\ell}\rangle_{n}$ and variance $v_{n}:=\langle(\hat{\ell}-\mu_{n})^{2}\rangle_{n}$; (ii) approximate the interface kernel $K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})$ by a Gaussian in $\ell$,
@@
with the conditional mean $\tilde{\mu}(\ell_{\mathrm{L}},\ell_{\mathrm{R}})$ and conditional variance $\tilde{v}(\ell_{\mathrm{L}},\ell_{\mathrm{R}})$ given above. Our exact numerics SM confirm (i) via the rapid decay of the Kolmogorov distance between the standardized $P_{n}$ and the standard normal distribution, while (ii) can be proved in an appropriate asymptotic regime.
@@
Here, the two contributions to $v_{n}$ come from the conditional variance in (7) and from the fluctuations of $\ell_{\mathrm{L}}$ and $\ell_{\mathrm{R}}$, respectively.
@@
which is independent of $d$. Therefore, $v_{n}/\mu_{n}$ rapidly converges to $2/3$ as $n\to\infty$ (the “$2/3$ law”), in agreement with our exact numerical iteration of the recursion SM .
Gaussian closure in d=2d=2.— In the marginal dimension $d=2$, the leading drift in (8) cancels; retaining the next-order terms predicts a logarithmic growth
@@
i.e., a weakly anomalous behavior. Note that the transport is extremely weak. The Gaussian closure still yields the 2/3 law, $v_{n}/\mu_{n}\to 2/3$. These conclusions, as well as the (slow) approach of $P_{n}$ to a Gaussian shape, are supported by numerics SM .
Remarkably, both the logarithmic behavior of the mean conductance and the $2/3$ law were obtained for the original Lorentz mirror model in $d=2$ by Nahum, Serna, Somoza, and Ortuño NahumSernaSomozaOrtuno2013 via a field-theoretic RG analysis and large-scale Monte Carlo simulations (up to $L=10^{6}$). See also ZiffKongCohen1991 ; OwczarekPrellberg for earlier numerical indications of logarithmic corrections.
@@
Numerical results for the original Lorentz mirror model in d=3d=3.— The rapid stabilization of $v_{n}/\mu_{n}$ to $2/3$ in the hierarchical model for $d\geq 2$, together with the RG analysis and large-scale simulations in NahumSernaSomozaOrtuno2013 for the original model in $d=2$, motivates viewing the variance-to-mean ratio of the conductance as a universal amplitude ratio, characterizing a universality class that contains the present hierarchical Lorentz mirror model. We shall now see that this expectation is justified, in a stronger sense than we initially anticipated.
Normal-transport scaling (1) in the original Lorentz mirror model for $d=3$ was first supported numerically in ChiffaudelLefevere . The proportionality constant (conductivity) $\kappa:=\lim_{L\uparrow\infty}\bar{\mu}(L)/L$ was derived in Lefevere2025 from a multiscale analysis under a closure hypothesis on trajectory correlations. This constant is nonuniversal and depends on the microscopic pairing rule.
@@
Discussion.— We introduced a hierarchical Lorentz mirror model that captures deterministic transport generated solely by quenched environmental randomness through random matching. For this model, we proved that the conductance exhibits normal-transport scaling, even though the microscopic dynamics is far from diffusive and many trajectories are closed, as in the original mirror model. More precisely, we established that $\mu_{n}\propto A_{n}/L_{n}$ in all dimensions $d\geq 3$. The marginal case $d=2$ remains open: the Gaussian approximation and numerics suggest that $\mu_{n}$ grows logarithmically with the system size, in agreement with the previous study of the non-hierarchical model NahumSernaSomozaOrtuno2013 .
A second outcome is the “$2/3$ law”: whenever $\mu_{n}$ grows with scale, the variance satisfies $v_{n}/\mu_{n}\to 2/3$. In the hierarchical model, this emerges from the Gaussian closure leading to the recursion (9), and is supported numerically. The same ratio 2/3 was already found in the non-hierarchical model with $d=2$ in NahumSernaSomozaOrtuno2013 . Moreover, we provide numerical evidence that the same ratio holds for the original Lorentz mirror model in $d=3$, and appears robust under variations of the local pairing rule. This motivates the conjecture that the “$2/3$ law” is universal within a broad class of models.
Our results suggest that $d=2$ is the unique marginal dimension of the hierarchical model, and we conjecture the same for the original model. A single injected trajectory may be viewed as a non-backtracking walk with strong memory (encoding the quenched random environment), suggesting a comparison with “true” (myopic) self-avoiding walks and related self-repelling processes, for which $d=2$ is also marginal and which become diffusive in $d\geq 3$ AmitParisiPeliti ; Toth1995 ; TothWerner1998 ; HorvathTothVeto2012 . At the same time, if crossings were produced by essentially independent diffusive trajectories, one would expect Poisson counting statistics, for which $v/\mu$ equals $1$ rather than approaching the sub-Poisson value $2/3$. This contrast suggests that the global current matching—rather than the scaling of an individual trajectory—is the mechanism behind normal transport with the $2/3$ law. A rigorous derivation of the $2/3$ law, and control of the marginal case $d=2$, remain important open problems.
@@ Acknowledgements.
We thank Hosho Katsura for pointing out that Ref. NahumSernaSomozaOrtuno2013 contains results closely related to the present study. R.L. thanks Gakushuin University for its hospitality during his stay in Tokyo, when the main part of this work was carried out. He also acknowledges the financial support from FJ-LMI IRL 2025 CNRS-The University of Tokyo, which made the visit possible. The present research is supported by JSPS Grants-in-Aid for Scientific Research Nos. 22K03474 (R.L. and H.T.) and 25K07171 (H.T.). We made extensive use of ChatGPT (versions 5.1 and 5.2, Plus/Pro) throughout this project for code development, assistance with proofs and calculations, reference searches, and drafting of the manuscript. In particular, the role of the identity (16) in the proof of Theorem End Matter was identified with the assistance of ChatGPT 5.2 Pro (although, interestingly, it never realized (16) until we pointed out that the quantity could be evaluated).
@@
In the main text, we also referred to numerical results for the three-dimensional Lorentz mirror model with the “orthogonal rule,” in which a particle is forced to turn at every visited vertex (equivalently, no edge is paired with its opposite). We expect that the models with the standard pairing rule and the orthogonal pairing rule belong to the same universality class in $d=3$ and exhibit the same macroscopic scaling behavior. This is in contrast with $d=2$, where the Lorentz mirror model with the orthogonal rule is known to belong to a different universality class than the standard model, namely that of the percolation critical point; see, e.g., ZiffKongCohen1991 ; NahumSernaSomozaOrtuno2013 .
@@
SM.2.6 Universal laws for higher cumulants
One may go beyond the Gaussian closure by allowing cubic and quartic corrections in the effective potential $\log P_{n}(\ell)$. A convenient way is to expand in the centered variable $\varphi:=(\ell-\mu_{n})/\sqrt{v_{n}}$, i.e., to supplement the Gaussian form with $\varphi^{3}$ and $\varphi^{4}$ terms and carry out a standard perturbative analysis of the recursion. This predicts that, for any $d\geq 2$, the third and fourth cumulants of the crossing number $\hat{\ell}$ obey the universal asymptotics (see [SM1])
$\kappa^{(3)}_{n}:=\langle(\hat{\ell}-\mu_{n})^{3}\rangle_{n}\simeq\frac{4}{15}\,\mu_{n},\qquad\kappa^{(4)}_{n}:=\langle(\hat{\ell}-\mu_{n})^{4}\rangle_{n}-3v_{n}^{2}\simeq-\frac{8}{105}\,\mu_{n},$ (SM.16)
as $n\uparrow\infty$. Equivalently, the ratios $\kappa^{(3)}_{n}/\mu_{n}$ and $\kappa^{(4)}_{n}/\mu_{n}$ approach the universal constants $4/15$ and 8/105-8/105, respectively, in the same sense as the $2/3$ law for $v_{n}/\mu_{n}$.
These higher-cumulant laws provide a stringent numerical test of universality for the (non-hierarchical) original Lorentz mirror model, especially in $d=3$. For $d=2$, the “$4/15$ law” was already obtained from a field-theoretic RG analysis and confirmed numerically in NahumSernaSomozaOrtuno2013 .
@@
$u_{n}=\sum_{\ell_{1},\ldots,\ell_{m}}\Bigl(1+\sum_{j=1}^{m}\ell_{j}\Bigr)^{-1}\prod_{j=1}^{m}P_{n}(\ell_{j}),$ (SM.17)
@@
$u_{n}=\sum_{y_{1},\ldots,y_{m}}\Bigl(\sum_{j=1}^{m}y_{j}^{-1}-(m-1)\Bigr)^{-1}\prod_{j=1}^{m}Q_{n}(y_{j}).$ (SM.18)
@@
$(y_{1}^{-1}+c)^{-1}=\frac{y_{1}}{1+cy_{1}}$ (SM.19)
is concave in $y_{1}\in(0,1]$, since its second derivative equals $2c(1+cy_{1})^{-3}\leq 0$. Applying Jensen’s inequality to the $y_{1}$-average in (SM.18) yields
$u_{n}\leq\sum_{y_{2},\ldots,y_{m}}\Bigl(\eta_{n}^{-1}+\sum_{j=2}^{m}y_{j}^{-1}-(m-1)\Bigr)^{-1}\prod_{j=2}^{m}Q_{n}(y_{j}).$ (SM.20)
@@
$u_{n}\leq\bigl(m\eta_{n}^{-1}-(m-1)\bigr)^{-1}=\frac{\eta_{n}}{m-(m-1)\eta_{n}}=\frac{\eta_{n}}{2^{d-1}-(2^{d-1}-1)\eta_{n}}.$ (SM.21)
@@
Substituting (SM.21) (with $n\mapsto n-1$) into (24) and using that $x\mapsto 2x-x^{2}$ is increasing on $[0,1]$, we obtain
$\eta_{n}\leq\frac{2m\,\eta_{n-1}-(2m-1)\eta_{n-1}^{2}}{\bigl(m-(m-1)\eta_{n-1}\bigr)^{2}}.$ (SM.22)
@@
$\qquad=-\frac{\eta^{2}\bigl(\eta(m-1)^{2}-m(m-2)\bigr)^{2}}{m^{2}\bigl(m-(m-1)\eta\bigr)^{2}}\leq 0.$ (SM.23)
Combining (SM.22) and (SM.23) yields the quadratic recursion
$\eta_{n}\leq\frac{2}{m}\eta_{n-1}+\Bigl(\frac{m-1}{m}\Bigr)^{2}\eta_{n-1}^{2}=2^{2-d}\eta_{n-1}+(1-2^{1-d})^{2}\eta_{n-1}^{2}.$ (SM.24)
@@
$x_{n}:=\alpha^{-n}\eta_{n}=\Bigl(\frac{m}{2}\Bigr)^{n}\eta_{n}=2^{(d-2)n}\eta_{n}.$ (SM.25)
Then (SM.24) implies
$x_{n}\leq x_{n-1}+\beta\,\alpha^{\,n-2}x_{n-1}^{2},\qquad\beta:=\Bigl(\frac{m-1}{m}\Bigr)^{2},$ (SM.26)
@@
$\frac{1}{x_{n}}\geq\frac{1}{x_{n-1}}-\beta\,\alpha^{\,n-2},$ (SM.27)
@@
$\frac{1}{x_{n}}\geq(A_{0}+1)-\beta\sum_{j=1}^{n}\alpha^{j-2}\geq(A_{0}+1)-\frac{\beta}{\alpha(1-\alpha)}=(A_{0}+1)-\frac{(m-1)^{2}}{2(m-2)}.$ (SM.28)
@@
$\eta_{n}^{-1}=\alpha^{-n}x_{n}^{-1}\geq\Bigl(A_{0}+1-\frac{(m-1)^{2}}{2(m-2)}\Bigr)\Bigl(\frac{m}{2}\Bigr)^{n}=\Bigl(A_{0}+1-\frac{(2^{d-1}-1)^{2}}{2(2^{d-1}-2)}\Bigr)2^{(d-2)n}.$ (SM.29)
@@
$\mu_{n}\geq\Biggl[1-\frac{1}{A_{0}}\biggl(\frac{(2^{d-1}-1)^{2}}{2(2^{d-1}-2)}-1\biggr)\Biggr]\frac{A_{n}}{L_{n}}-1,$ (SM.30)
@@
The present section provides additional information on our numerical work. The codes are archived on Zenodo (DOI: 10.5281/zenodo.18622214). We focus on the hierarchical Lorentz mirror model, where the recursion (2) can be iterated essentially exactly (i.e., without Monte Carlo sampling), and we briefly discuss the original (non-hierarchical) Lorentz mirror model in the final subsection.
@@
$P_{n-1}^{*m}(\ell^{\prime}):=\mathop{\sum_{\ell_{1},\ldots,\ell_{m}}}_{(\sum_{j=1}^{m}\ell_{j}=\ell^{\prime})}\prod_{j=1}^{m}P_{n-1}(\ell_{j}),$ (SM.31)
@@
$P_{n}(\ell)=\sum_{\ell_{\mathrm{L}},\ell_{\mathrm{R}}}K(\ell\,|\,\ell_{\mathrm{L}},\ell_{\mathrm{R}})\,P_{n-1}^{*m}(\ell_{\mathrm{L}})\,P_{n-1}^{*m}(\ell_{\mathrm{R}}).$ (SM.32)
@@
$\mu_{n}\simeq a+b\,n\quad\text{(as $n\uparrow\infty$)},\qquad b\simeq 0.08346,$ (SM.33)